Hi my name is Lissa. This week in my college algebra 2 class we are working on Rational Exponents and Radical Expressions. This stuff is so confusing to me, I need help with this problem please.
Simplify; express answer with only positive exponents.
(a^-1/4 b^2/3)(a^1/3 b^3/4)

Answer by stanbon(75887)

Simplify; express answer with only positive exponents.
(a^-1/4 b^2/3)(a^1/3 b^3/4)
-----------------------------
Check the "a" factors: a^(-1/4)*a(1/3) = a^(1/3 - 1/4) = a^(1/12)
-------------------
Check the "b" factors:b^(2/3)*b(3/4)=b^(2/3 + 4/4)=b^(20/6)=b^(10/3)
------------------------
Ans: a^(1/12)*b^(10/3)
==========================
Cheers,
Stan H.